# The role of prompt cusps in driving the core collapse of SIDM halos

### v1

@@ II.1 Prompt cusp density profiles and SIDM model
We simulate halos that could have formed in a universe with the particle physics model used by [27] with a kinematic decoupling temperature $T_{\rm{kd}}=1.46\,{\rm keV}$. This model exhibits a cutoff in the matter power spectrum comparable to that of a $6.5\,{\rm keV}$ thermal relic warm DM model, with a suppression of small-scale structure that manifests on scales between $10^{7}\text{--}10^{8}{\,\rm M_{\odot}}$. Due to the suppression of small-scale power, low-mass halos become both less abundant and less concentrated, with the latter effect being of particular relevance for the gravothermal evolution of SIDM halos, given that the characteristic timescale for core collapse scales as $c^{-7/2}$, where $c$ is the concentration of an NFW profile [30].
We aim to understand how a prompt cusp at the center of a low-mass halo impacts the time evolution of the halo profile. For what follows, we assume a halo of mass $10^{7}{\,\rm M_{\odot}}$ with a concentration $c=4.74$ at redshift $z=3$, comparable to some of the lowest-mass structures expected to form for the corresponding matter power spectrum. The concentration is defined to be $c=r_{200}/r_{-2}$, where $r_{200}$ is the virial radius and $r_{-2}$ is the radius where the logarithmic density profile slope crosses 2-2. Using the model presented by  [18], we inject prompt cusps into the centers of these halos. The composite PC+NFW halo profile is given by [18]
@@
We assume an SIDM cross section
@@
with $\sigma_{0}/m=147\ \rm{cm^{2}}\rm{g^{-1}}$ and $w=24\ \rm{km}\ \rm{s^{-1}}$. This cross section corresponds to the $T_{\rm{kd}}=1.46\rm{keV}$ SIDM model, which exhibits a small-scale cutoff in the matter power spectrum expected to give rise to halos comparable to the three cases described above. This parametric model for the scattering cross section closely matches the cross section that results from a Yukawa potential in the Born approximation. Following [31, 32, 29], we approximate halos evolved under such a velocity-dependent cross section (VDCS) by an effective velocity-independent cross section (VICS) with a strength given by [31, 32]
@@ II.2 N-body simulation details
All halos are initialized using stable configurations derived in the CDM model, with adjustments made to account for the softened gravitational field, as detailed in Appendix A. The initial conditions (ICs) of our halos are generated following the parametric density profile detailed in Equation 1 within the virial radius ($r_{200}$) and a modified exponential cut-off outside $r_{200}$ [40, 41]. The profiles take the analytical forms of
$\rho_{r\leq r_{200}}(r)$ $=\rho(r\!\mid\!r_{\rm{s}},\rho_{\rm{s}},y)$ (6)
$\rho_{r>r_{200}}(r)$ $=\rho_{200}\left(\frac{r}{r_{200}}\right)^{\epsilon_{\rm d}}\exp\left(-\frac{r-r_{200}}{r_{\rm d}}\right),$ (7)
with $\rho_{200}=\rho(r_{200}\!\mid\!r_{\rm{s}},\rho_{\rm{s}},y)$. $r_{\rm d}$ is the decay scale radius, taken to be $r_{200}$ for simplicity, while $\epsilon_{\rm d}$ is the exponential decay index, chosen so that the continuity of the logarithmic slope of the density profile is preserved
@@
Figure 2 shows the initial density, enclosed mass, and one-dimensional velocity dispersion ratios between the prompt-cusp halos and control NFW configurations. Each prompt-cusp halo exhibits a higher central density than the NFW profile, owing to the presence of the PS, and also an enhanced density beyond $r_{200}$. In this outer region, the density profile of the extreme-cusp halo deviates from those of the low-cusp and control halos by $10\text{--}15\%$, and from the high-cusp halo by $\sim 5\%$, corresponding to enclosed-mass deviations of $\sim 5\%$ and $\sim 2\%$, respectively. These differences arise directly from the chosen parameterization of the halo density profile, as well as from the exponential cutoff applied to the outer regions. The increased mass at large radii also produces a corresponding increase in the velocity dispersion in those regions. Although these deviations are relatively small, they may nonetheless affect the subsequent collapse times by altering the temperature gradient, and consequently the efficiency of heat transport.
@@ III.1 The general evolution of halo structures
Similar to what has been observed in other studies of elastic DM self-interactions [e.g. 39, 32, 29], our halos undergo an initial core-formation phase, characterized by the development of an isothermal, constant-density core, followed by gravothermal core collapse [42]. Figure 3 shows the evolution of the high-cusp and control NFW halos during the core-formation phase. Due to being initially much denser, owing to the prompt cusp, the high-cusp halo develops a core at a noticeably slower pace than the NFW reference. This trend is evident in both the density and velocity dispersion profiles. Nevertheless, after $T\sim 3\,{\rm Gyr}$, the two halos begin to display comparable internal kinetic structures, even though their density profiles still differ by several tens of percent in the central regions. To quantify the evolution of SIDM halos more precisely, we characterize the SIDM halo cores using the core density ($\rho_{\rm{c}}$), the core half-density radius ($r_{\rho/2}$, defined as the radius at which the density drops to half the value at the core), and the core one-dimensional velocity dispersion ($\sigma_{\rm{1D,c}}$), following the procedures detailed in [29]. Specifically, $\sigma_{\rm{1D,c}}$ is measured directly from the one-dimensional velocity dispersion profiles of the halos using an iterative hypothesis testing procedure (following Section III.C of [29]). $\rho_{\rm{c}}$ and $r_{\rho/2}$ are obtained by fitting the density profile using the function form
@@ III.3 Late-time evolution & core-collapse time
Table 3: The relative collapse times of the prompt cusp halos with respect to the reference NFW halo. The empirical collapse time $\tau_{\rm 100}$ is defined as the epoch when the core density reaches $100$ times the effective scale density $\bar{\rho}_{\rm{s}}$. The ratios of the effective collapse timescale $\bar{\tau}\propto 1/\bar{\rho}_{\rm{s}}\bar{\sigma}_{\rm{s}}$ are also shown. Such ratios, however, do not appear to serve as reliable predictors of the collapse time in prompt cusp halos.
@@
As discussed in Section II.2, excess mass residing beyond $r_{200}$ observed in Figure 2 can plausibly decelerate the collapse time of halos: this additional material increases the temperature in the outskirts, suppressing the outward heat transport and slowing down the core collapse. We see this trend for NFW profiles as well, shown in Figure 9 (see Appendix B), where the collapse of the control NFW halo is accelerated by the deficit of material in the outer regions compared to a more extended configuration. However, not every halo with excess material in the outskirts exhibits decelerated core collapse; the low and high-cusp halos both collapse slightly faster than the NFW control. On the other hand, the extreme-cusp halo collapses more slowly.
From these observations, we infer that there may be multiple competing factors that govern the overall collapse timescale. Among them is the increase in central density, which acts to accelerate core collapse. The second likely contributing factor is the influence of the outer boundary, which can either impede or enhance the collapse depending on the thermal state of the outer region. The same observations made in this section for halos evolved under the VICS model also apply to halos evolved under VDCS (see Appendix D). The slight deviations between the two cases can be attributed to the additional complexity in the collapse times introduced by the velocity dependence.
@@ Acknowledgements.
We thank Francis-Yan Cyr-Racine and Hai-bo Yu for useful discussions. The simulations in this paper were conducted on the Engaging cluster at Massachusetts Institute of Technology (MIT) operated by the MIT Office of Research Computing and Data. DG acknowledges support for this work provided by the Brinson Foundation through a Brinson Prize Fellowship grant. XS acknowledges the support from the National Aeronautics and Space Administration (NASA) theory grant JWST-AR-04814. Support for OZ was provided by Harvard University through the Institute for Theory and Computation Fellowship.
@@ A.1 Eddington’s distribution function
$f\left(\xi\right)=\frac{1}{\sqrt{8}\pi^{2}}\int_{0}^{\xi}{\rm d}\psi\left[\frac{1}{\sqrt{\xi-\psi}}\frac{{\rm d}^{2}\rho}{{\rm d}\psi^{2}}+\frac{1}{\sqrt{\xi}}\frac{{\rm d}\rho}{{\rm d}\psi}\Bigr\rvert_{\psi=0}\right].$ (14)
@@ A.2 ICs in the context of softened gravity
Typically, the Eddington distributions are constructed with the Newtonian gravitational potential. However, under the softened gravity schemes employed in simulations, even without DM self-interaction, halos sampled using such distributions would become unstable, leading to the formation of small cores roughly the size of the softening kernel [e.g. 28, 44]. This is shown in Figure 8, where we evolved the SIDM-control halo in the CDM model under a large softening length, $10$ times the recommended value in [38].
@@
$W_{\rm{S}}(x;h)=\frac{8}{\pi h^{3}}\begin{cases}1-6\left(\frac{x}{h}\right)^{2}+6\left(\frac{x}{h}\right)^{3},&0\leq\frac{x}{h}<\frac{1}{2}\\ 2\left(1-\frac{x}{h}\right)^{3},&\frac{1}{2}\leq\frac{x}{h}<1\\ 0,&\frac{x}{h}\geq 1\end{cases}$ (18)
@@
$\mathcal{I}(u_{1},u_{2})=\begin{cases}A(u_{2})-A(u_{1}),&u_{2}<\frac{1}{2}\\ B(u_{2})-B(\frac{1}{2})+A(\frac{1}{2})-A(u_{1}),&u_{1}\leq\frac{1}{2}\leq u_{2}\\ B(u_{2})-B(u_{1}),&u_{1}\geq\frac{1}{2}\\ 0,&u_{1}\geq u_{2}\end{cases}$ (20)
Here, the final case occurs when $r$ is outside of the kernel width, i.e., $\left|r-R\right|\geq h$. The two auxiliary equations $A(u)$ and $B(u)$ are defined as
$A(u)$ $=\frac{1}{2}u^{2}-\frac{3}{2}u^{4}+\frac{6}{5}u^{5},$ (21)
$B(u)$ $=u^{2}-2u^{3}+\frac{3}{2}u^{4}-\frac{2}{5}u^{5}.$ (22)
In simulations that employ such a spline softening approach, such as those in AREPO and GADGET, the gravitational field resulting from an isotropic density profile $\rho(r)$ would be equivalent to the Newtonian field generated by the convoluted density profile
$\rho_{\rm{S}}(r)=\int_{0}^{\infty}\frac{16r^{\prime}\rho(r^{\prime})\,{\rm{d}}r^{\prime}}{hr}\\ \mathcal{I}\left(\left|r-r^{\prime}\right|/h,\,\max{\left(1,\,\left(r+r^{\prime}\right)/h\right)}\right).$ (23)
The gravitational potential would be calculated as
$\Phi_{\rm{S}}(r)=-G\left[\frac{1}{r}\int_{0}^{r}4\pi{r^{\prime}}^{2}\rho_{\rm{S}}(r^{\prime})\,{\rm{d}}r^{\prime}+\right.\\ \left.\int_{r}^{\infty}4\pi r^{\prime}\rho_{\rm{S}}(r^{\prime})\,{\rm{d}}r^{\prime}\right].$ (24)
@@ Appendix B Accelerated evolution due to outer halo stripping
Figure 9 compares the core evolution track of the SIDM-control halo in the VICS model with the approximate self-similar track from [35], obtained using the predicted timescale of $\tau=43.13\,{\rm Gyr}$. For comparison, we also include an “extended” version of the SIDM-control halo, initialized with an NFW profile (Equation 6) out to $5,r_{200}$, rather than switching to the exponential cutoff at $r_{200}$. This extended halo effectively represents a more concentrated configuration in the cosmological context. Initially, the core evolution of both halos follows the self-similar track closely. However, as the systems enter the core-collapse phase, deviations emerge. While the extended halo continues to follow the self-similar reference, the SIDM-control halo exhibits an approximately $10\%$ faster collapse. Moreover, we observe the transition index of the SIDM-control halo reaches lower values, indicating a relative deficiency of material in the outer regions, which weakens the contribution from the $\rho\propto r^{-3}$ regime. This deficit also results in a cooler outer region, promoting faster outward heat transport and thereby accelerating the gravothermal collapse. The fact that these deviations arise only in the later stages can be attributed to the varying collision timescales across the halo: at larger radii, where both the density and velocity dispersion are lower, collisions occur less frequently, delaying their influence on the evolution of the core. This is consistent with the effects of tidal stripping observed in cosmological simulations [e.g. 48].
@@ Appendix C The effective NFW parameters of prompt cusp halos
Appendix D Halo comparison in the VDCS model

### v2

# Role of prompt cusps in driving the core collapse of self-interacting dark matter halos

@@ II.1 Prompt cusp density profiles and SIDM model
Building on [6], we consider the formation of PC in the early Universe in the context of SIDM. Although [6] did not explicitly consider the effects of SIDM on PC formation, in order to study the late-time impact of PC on halo evolution, we assume that the formation process proceeds in much the same way as it would in warm dark matter (WDM). This is a reasonable assumption as in the case of WDM, the initial collapses of PCs begin around $z\sim 10\text{--}20$, after which it takes roughly $25\%$ of the Hubble time for them to fully form [6]. At $z=10$, this corresponds to $\sim 0.2\,{\rm Gyr}$, which is much shorter than the timescale we later consider for self-interactions of DM. Therefore, we believe there is no significant difference in the PC formation process between WDM and SIDM with small-scale structure suppression.
We simulate halos that could have formed in a universe with the particle physics model used by [27] with a kinematic decoupling temperature $T_{\rm{kd}}=1.46\,{\rm keV}$. This model exhibits a cutoff in the matter power spectrum comparable to that of a $6.5\,{\rm keV}$ thermal relic WDM model, with a suppression of small-scale structure that manifests on scales between $10^{7}\text{--}10^{8}{\,\rm M_{\odot}}$. We aim to understand how a prompt cusp at the center of a low-mass halo impacts the time evolution of the halo profile. For what follows, we assume a halo of virial mass $M_{200}=10^{7}{\,\rm M_{\odot}}$ with a concentration $c=4.74$ at redshift $z=3$, comparable to some of the lowest-mass structures expected to form for the corresponding matter power spectrum. The concentration is defined to be $c=r_{200}/r_{-2}$, where $r_{200}$ is the virial radius and $r_{-2}$ is the radius where the logarithmic density profile slope crosses 2-2. We note that here, the assumption that self-interactions do not have a significant impact on the halo evolution prior to $z=3$ is adopted. This may not be entirely correct, as seen later in the evolution of the DM halos, and represents a common limitation of SIDM simulations that start from NFW-like ICs. Using the model presented by  [18], we inject prompt cusps into the centers of these halos. The composite PC+NFW halo profile is given by [18]
@@
We assume an isotropic SIDM cross section
@@
with $\sigma_{0}/m=147\ \rm{cm^{2}}\rm{g^{-1}}$ and $w=24\ \rm{km}\ \rm{s^{-1}}$. This cross section corresponds to the $T_{\rm{kd}}=1.46\rm{keV}$ SIDM model, which exhibits a small-scale cutoff in the matter power spectrum expected to give rise to halos comparable to the three cases described above. The values of this cross section place the halos in the long mean-free-path (LMFP) regime throughout their evolution. This parametric model for the scattering cross section closely matches the cross section that results from a Yukawa potential in the Born approximation. In the case of DM halos evolved with self-interactions from NFW ICs, the evolution can be approximated using an effective cross section and collapse timescale. We examine these scale parameters to determine whether the same description remains valid for halos embedded with PCs. Following [30, 31, 29], we approximate halos evolved under such a velocity-dependent cross section (VDCS) by an effective velocity-independent cross section (VICS) with a strength given by [30, 31]
@@ II.2 N-body simulation details
All halos are initialized using stable configurations derived in the CDM model, with adjustments made to account for the softened gravitational field, as detailed in Appendix A. The initial conditions (ICs) of our halos are generated following the parametric density profile detailed in Equation 1 within the virial radius ($r_{200}$) and a modified exponential cut-off outside $r_{200}$ [39, 40]. The profiles take the analytical forms of {align} ρ_r ≤r_200 (r) = ρ(r ∣ r_s, ρ_s, y)
ρ_r ¿ r_200 (r) = ρ_200(rr200)^ϵ_d exp(-r-r200rd), with $\rho_{200}=\rho(r_{200}\!\mid\!r_{\rm{s}},\rho_{\rm{s}},y)$. $r_{\rm d}$ is the decay scale radius, taken to be $r_{200}$ for simplicity, while $\epsilon_{\rm d}$ is the exponential decay index, chosen so that the continuity of the logarithmic slope of the density profile is preserved
@@
Figure 2 shows the initial density, enclosed mass, and one-dimensional velocity dispersion ratios between the prompt-cusp halos and control NFW configurations. Each prompt-cusp halo exhibits a higher central density than the NFW profile, owing to the presence of the PC, and also an enhanced density beyond $r_{200}$. In this outer region, the density profile of the extreme-cusp halo deviates from those of the low-cusp and control halos by $10\text{--}15\%$, and from the high-cusp halo by $\sim 5\%$, corresponding to enclosed-mass deviations of $\sim 5\%$ and $\sim 2\%$, respectively. These differences arise directly from the chosen parameterization of the halo density profile, as well as from the exponential cutoff applied to the outer regions. The increased mass at large radii also produces a corresponding increase in the velocity dispersion in those regions. Although these deviations are relatively small, they may nonetheless affect the subsequent collapse times by altering the temperature gradient, and consequently the efficiency of heat transport.
@@ III.1 The general evolution of halo structures
Similar to what has been observed in other studies of elastic DM self-interactions [e.g. 38, 31, 29], our halos undergo an initial core-formation phase, characterized by the development of an isothermal, constant-density core, followed by gravothermal core collapse [41]. Figure 3 shows the evolution of the high-cusp and control NFW halos during the core-formation phase. Due to being initially much denser, owing to the prompt cusp, the high-cusp halo develops a core at a noticeably slower pace than the NFW reference. This trend is evident in both the density and velocity dispersion profiles. Nevertheless, after $T\sim 3\,{\rm Gyr}$, the two halos begin to display comparable internal kinetic structures, even though their density profiles still differ by several tens of percent in the central regions. To quantify the evolution of SIDM halos more precisely, we characterize the SIDM halo cores using the core density ($\rho_{\rm{c}}$), the core half-density radius ($r_{\rho/2}$, defined as the radius at which the density drops to half the value at the core), and the core one-dimensional velocity dispersion ($\sigma_{\rm{1D,c}}$), following the procedures detailed in [29]. Specifically, $\sigma_{\rm{1D,c}}$ is measured directly from the one-dimensional velocity dispersion profiles of the halos using an iterative hypothesis testing procedure (following Section III.C of [29]). $\rho_{\rm{c}}$ and $r_{\rho/2}$ are obtained by fitting the density profile using the function form
@@ III.3 Late-time evolution & core-collapse time
Table 3: The relative collapse times of the prompt-cusp halos with respect to the reference NFW halo. The empirical collapse time $\tau_{\rm 100}$ is defined as the epoch at which the core density reaches $100$ times the effective scale density $\bar{\rho}_{\rm{s}}$. The ratios of the effective collapse timescale $\bar{\tau}\propto 1/\bar{\rho}_{\rm{s}}\bar{\sigma}_{\rm{s}}$ are also shown. Such ratios, however, do not appear to serve as reliable predictors of the collapse time in prompt cusp halos.
@@
As discussed in Section II.2, excess mass residing beyond $r_{200}$ observed in Figure 2 can plausibly decelerate the collapse time of halos: this additional material increases the temperature in the outskirts, decreasing the gradient and suppressing the outward heat transport, which slows down the core collapse. We see this trend for NFW profiles as well, shown in Figure 9 (see Appendix B), where the collapse of the control NFW halo is accelerated by the deficit of material in the outer regions compared to a more extended configuration. However, not every halo with excess material in the outskirts exhibits decelerated core collapse; the low and high-cusp halos both collapse slightly faster than the NFW control. On the other hand, the extreme-cusp halo collapses more slowly.
From these observations, we infer that there may be multiple competing factors that govern the overall collapse timescale. Among them is the increase in central density, which acts to accelerate core collapse. The second likely contributing factor is the influence of the outer boundary, which can either impede or enhance the collapse depending on the thermal state of the outer region. The same observations made in this section for halos evolved under the VICS model also apply to halos evolved under VDCS (see Appendix E). The slight deviations between the two cases can be attributed to the additional complexity in the collapse times introduced by the velocity dependence.
@@ Acknowledgements.
We thank Francis-Yan Cyr-Racine, Hai-bo Yu, and Moritz Fischer for useful discussions. The simulations in this paper were conducted on the Engaging cluster at Massachusetts Institute of Technology (MIT) operated by the MIT Office of Research Computing and Data. DG acknowledges support for this work provided by the Brinson Foundation through a Brinson Prize Fellowship grant. XS acknowledges the support from the National Aeronautics and Space Administration (NASA) theory grant JWST-AR-04814. Support for OZ was provided by Harvard University through the Institute for Theory and Computation Fellowship.
@@ A.1 Eddington’s distribution function
[…]
@@ A.2 ICs in the context of softened gravity
Typically, the Eddington distributions are constructed with the Newtonian gravitational potential. However, under the softened gravity schemes employed in simulations, even without DM self-interaction, halos sampled using such distributions would become unstable, leading to the formation of small cores roughly the size of the softening kernel [e.g. 28, 43]. This is shown in Figure 8, where we evolved the SIDM-control halo in the CDM model under a large softening length, $10$ times the recommended value in [37].
@@
$W_{\rm{S}}(x;h)=\frac{8}{\pi h^{3}}\cases{1}-6\left(\frac{x}{h}\right)^{2}+6\left(\frac{x}{h}\right)^{3},&0\leq\frac{x}{h}<\frac{1}{2}\\ 2\left(1-\frac{x}{h}\right)^{3},&\frac{1}{2}\leq\frac{x}{h}<1\\ 0,&\frac{x}{h}\geq 1$ (16)
@@
$\mathcal{I}(u_{1},u_{2})=\cases{A}(u_{2})-A(u_{1}),&u_{2}<\frac{1}{2}\\ B(u_{2})-B(\frac{1}{2})+A(\frac{1}{2})-A(u_{1}),&u_{1}\leq\frac{1}{2}\leq u_{2}\\ B(u_{2})-B(u_{1}),&u_{1}\geq\frac{1}{2}\\ 0,&u_{1}\geq u_{2}$ (18)
Here, the final case occurs when $r$ is outside of the kernel width, i.e., $\left|r-R\right|\geq h$. The two auxiliary equations $A(u)$ and $B(u)$ are defined as {align} A (u) = 12 u^2 - 32 u^4 + 65 u^5 ,
B (u) = u^2 - 2 u^3 + 32 u^4 - 25 u^5.
In simulations that employ such a spline softening approach, such as those in AREPO and GADGET, the gravitational field resulting from an isotropic density profile $\rho(r)$ would be equivalent to the Newtonian field generated by the convoluted density profile {multline} ρ_S (r) = ∫_0^∞16 rρ(r)   drh r
I ( — r - r^′— / h,   max(1,   (r + r^′) / h) ). The gravitational potential would be calculated as {multline} Φ_S (r) = -G [ 1r ∫_0^r 4πr^′^2 ρ_S(r^′)   d r^′+
∫_r^∞ 4πr^′ρ_S(r^′)   dr^′].
@@ Appendix B Accelerated evolution due to outer halo stripping
Figure 9 compares the core evolution track of the SIDM-control halo in the VICS model with the approximate self-similar track from [34], obtained using the predicted timescale of $\tau=43.13\,{\rm Gyr}$. For comparison, we also include an “extended” version of the SIDM-control halo, initialized with an NFW profile (Equation 2) out to $5,r_{200}$, rather than switching to the exponential cutoff at $r_{200}$. This extended halo effectively represents a more concentrated configuration in the cosmological context. Initially, the core evolution of both halos follows the self-similar track closely. However, as the systems enter the core-collapse phase, deviations emerge. While the extended halo continues to follow the self-similar reference, the SIDM-control halo exhibits an approximately $10\%$ faster collapse. Moreover, we observe the transition index of the SIDM-control halo reaches lower values, indicating a relative deficiency of material in the outer regions, which weakens the contribution from the $\rho\propto r^{-3}$ regime. This deficit also results in a cooler outer region, promoting faster outward heat transport and thereby accelerating the gravothermal collapse. The fact that these deviations arise only in the later stages can be attributed to the varying collision timescales across the halo: at larger radii, where both the density and velocity dispersion are lower, collisions occur less frequently, delaying their influence on the evolution of the core. This is consistent with the effects of tidal stripping observed in cosmological simulations [e.g. 47].
@@ Appendix C The effective NFW parameters of prompt cusp halos
Appendix D Supplemental prompt-cusp halos evolution
Figure 11: Core density evolution of prompt-cusp halos with varying levels of PC prominence, shown in physical units (main panel) and in terms of the effective scale density (sub-panel). As observed in Figures 4 and 5, the collapse time depends non-trivially on both the effective concentration and the configuration of the outer regions. The smooth variation of the collapse time as a function of $y$ further indicates that this effect is physical rather than numerical.
Figure 12: Comparison between the effective NFW concentrations calculated following Appendix C and those measured following Section III.1 for the halos presented in Figure 11. The dashed line indicates the case of perfect agreement, which is nearly achieved.
   $r_{\rm{s}}$    $\log\rho_{\rm{s}}$    $y$
   $\rm kpc$]    ${\rm M}_{\odot}\,{\rm kpc}^{-3}$]
   0.387    7.186    0.310
   0.405    7.133    0.378
   0.427    7.072    0.445
   0.511    6.867    0.635
   0.576    6.726    0.756
   0.643    6.595    0.878
Table 4: Halo configurations for the additional prompt-cusp halos evolved in Appendix D. The numerical setups of these runs are similar to those adopted in the VICS run described in Section II.2.
In order to further support the points made in Section III.1 and Appendix C, we evolve six additional halos with configurations detailed in Table 4. The parameters are chosen such that the PC prominence of these halos lies between those of the SIDM-low and SIDM-extreme cases. Figure 11 shows the evolution of the core density for these halos, alongside the evolution of the halos presented in the main text. As expected from the dual influence of the increased concentration in the inner regions and the elevated temperature of the outer regions, the collapse time varies smoothly but non-trivially. This supports the argument made in Section III.3. Figure 12 shows the effective NFW concentrations measured directly from the simulations, as well as those calculated following the predictions of Appendix C. These values are consistent and demonstrate that the mass within $\simeq 0.65\,r_{-2}$ can be used to predict the behavior of halos with PC (and potentially other configurations) during the core formation phase.
Appendix E Halo comparison in the VDCS model